OpenStudy (anonymous):

What is the length of the longest side of a triangle that has vertices at (-5, 2) (1,-6) and (1,2)?

OpenStudy (anonymous):

Use the distance formula

OpenStudy (anonymous):

Distance formula?

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

ok its \[d=\sqrt{(x _{1}-x _{2})^2+(y _{1}-y _{2})^2} \]

OpenStudy (mathlegend):

I would also suggest labeling your coordinates

OpenStudy (anonymous):

MathLegend ur not a math legend at all

OpenStudy (mathlegend):

A= (-5, 2) B = (1,-6) C = (1,2)

OpenStudy (mathlegend):

So when using the distance formula... you can solve for side AB first.

OpenStudy (anonymous):

Thank you.

OpenStudy (mathlegend):

So using that formula, can you label the coordinates as (x1, y1) & (x2,y2)

OpenStudy (hba):



OpenStudy (mathlegend):

A= (-5, 2) B = (1,-6) So lets solve for side AB first... so, "A" comes first right? So let that be your x1 & y1 Then, "B" can be your x2 & y2

OpenStudy (mathlegend):

Do you understand so far @ValentinaT ?

OpenStudy (mathlegend):

All I did was label them so that we can plug it into that distance formula.

OpenStudy (mathlegend):

@ValentinaT let me know when you get back so we can work this out together. :)

OpenStudy (anonymous):

Yeah, I'm getting it, thank you.

OpenStudy (anonymous):

\[\frac{ -6 - 2 }{ -5 - 1 } = \frac{ -8 }{ -6}\]

OpenStudy (mathlegend):

So for AB (1+5)^2+(-6-2)^2

OpenStudy (mathlegend):

I'm sorry the formula is actually x2-x1 and y2-y1 the above poster just mixed up the two.

OpenStudy (anonymous):

Okay.

OpenStudy (mathlegend):

(1+5)^2+(-6-2)^2 (6)^2+(-8)^2 36+64

OpenStudy (mathlegend):

36+64 = 100

OpenStudy (mathlegend):

\[\sqrt{100}\]

OpenStudy (anonymous):

10

OpenStudy (mathlegend):

Now, we need the square root... because if you notice that entire formula had the square root symbol over it

OpenStudy (mathlegend):

Good.

OpenStudy (mathlegend):

So side AB = 10

OpenStudy (mathlegend):

So that is one side. So lets go for side BC

OpenStudy (mathlegend):

B = (1,-6) x1 y1 C = (1,2) x2 y2

OpenStudy (mathlegend):

@ValentinaT do you feel comfortable trying it out on your own? Tell me what you get and I'll check to see if your answer for side BC is correct.

OpenStudy (anonymous):

\[\frac{ 2 - -6 }{ 1 - 1 } = \frac{ 8 }{ 0 }\]

OpenStudy (mathlegend):

Remember we are not trying to find a slope.

OpenStudy (mathlegend):

We are looking for the distance between the vertices.

OpenStudy (mathlegend):

(1-1)^2 + (2+6)^2

OpenStudy (mathlegend):

To get that all I did was plug it into the formula. (x2-x1)^2+(y2-y1)^2

OpenStudy (mathlegend):

(1-1)^2 + (2+6)^2 Try solving that.

OpenStudy (anonymous):

Okay, 0 + 64?

OpenStudy (mathlegend):

Good

OpenStudy (mathlegend):

0 + 64 = 64 \[\sqrt{64}\]

OpenStudy (anonymous):

So, 8 as the square root.

OpenStudy (mathlegend):

Yes, so side BC = 8

OpenStudy (mathlegend):

Now, try side AC

OpenStudy (mathlegend):

(x2-x1)^2+(y2-y1)^2

OpenStudy (anonymous):

A = -5, 2 = x1, y1 C = 1, 2 = x2, y2 \[\frac{ 1 - -5^2}{ 2 - 2^2 } = \frac{ 6 }{ 1 } \] so 36?

OpenStudy (mathlegend):

(x2-x1)^2+(y2-y1)^2 (1+5)^2 + (2-2)^2 Do you understand this step?

OpenStudy (anonymous):

36 + 0 Yeah I get it, I just couldn't figure out how to write it.

OpenStudy (mathlegend):

Good so if you took the square root of 36 \[\sqrt{36}\]

OpenStudy (mathlegend):

So now, we know the length of the longest side is AB

OpenStudy (anonymous):

Okay, thank you!